STATISTICAL ERROR PROPAGATION IN THE COMPUTATION OF RKR POTENTIAL CURVES, SPEC-TROSCOPIC CONSTANTS, AND FRANCK-CONDON FACTORS: THE D STATE OF I_{2}

J. TELLINGHUISEN, Department of Chemistry, Vanderbilt University, Nashville, TN 37235.

Existing spectroscopic data for the D ion-pair state of I_{2} are reanalyzed with the primary goal of improving and extending to higher v the rotational constants B_{v}. The lack of rotational data for the v = 130-200 region makes the structure of this data set unusual, and has prompted a thorough study of the propagated statistical error in the directly fitted spectroscopic constants, T_{v} and B_{v}, and in the RKR potential curve calculated therefrom. For the first time, such error propagation computations have been extended to properties computed by numerical solution of the Schroedinger equation for the state in question: the quantum vibrational energy G_{v} and rotational constant B_{v}, centrifugal distortion constants D_{v} and H_{v}, and Franck-Condon factors for the D − X transition.